# Uni-MuMER: Unified Multi-Task Fine-Tuning of Vision-Language Model for Handwritten Mathematical Expression Recognition

### v2

@@ D.2 Results on Printed Mathematical Expressions
An important aspect of our approach is whether a model fine-tuned primarily for handwritten expressions also performs well on printed mathematical expressions. To test this, we evaluate on the Im2LaTeXv2 dataset, a dataset of printed formulas rendered from . Tab. 14 summarizes these results using common metrics from the Im2LaTeXv2: Edit Score, BLEU-4, and ExpRate. We also include ExpRate@CDM for completeness, especially since our models output that could be visually equivalent even if not exactly matching. Our Vanilla baseline performed slightly better than the previous SOTA. After completing three data-driven tasks, Uni-MuMER further reached an ExpRate of 88.52. We also check our Uni-MuMER, which is primarily optimized for handwriting. Interestingly, its performance on purely printed data is a bit lower than Uni-MuMER, perhaps because training on such a broad mix causes a slight domain shift. Nonetheless, all our models outperform MathNet.
Table 14: Comparison of model performance on the im2latexv2 test set. We compare our models to prior Im2LaTeX systems. Metrics shown are Edit Score, BLEU-4, ExpRate (Exp), and ExpRate@CDM (@CDM)
Model Train Dataset Edit [%] Bleu-4 [%] Exp [%] @CDM [%]
WYGIWYS im2latex-100k 37.2 23.9 0.0 -
I2l-strips im2latex-140k 75.9 65.9 10.3 -
I2l-nopool im2latex-140k 76.0 66.4 10.4 -
MathNet im2latexv2 97.2 96.8 83.9 -
@@
Uni-MuMER our 1.6M samples 98.98 98.10 84.32 88.99

### v3

@@ D.2 Results on Printed Mathematical Expressions
[…]
Tab. 24 summarizes these results using common metrics from the Im2LaTeXv2: Edit Score, BLEU-4, and ExpRate. We also include ExpRate@CDM for completeness, especially since our models output that could be visually equivalent even if not exactly matching. Our Vanilla baseline performed slightly better than the previous SOTA. After completing three data-driven tasks, Uni-MuMER further reached an ExpRate of 88.52. We also check our Uni-MuMER, which is primarily optimized for handwriting. Interestingly, its performance on purely printed data is a bit lower than Uni-MuMER, perhaps because training on such a broad mix causes a slight domain shift. Nonetheless, all our models outperform MathNet [44].
Table 24: Comparison of model performance on the im2latexv2 test set. We compare our models to prior Im2LaTeX systems. Metrics shown are Edit Score, BLEU-4, ExpRate, and ExpRate@CDM. All metrics are percentages (%), and higher values $\uparrow\,$ indicate better performance. The lower half reports results on the original, unnormalized test set. Missing values indicated by ’–’.
Model Train Dataset Edit Bleu-4 ExpRate ExpRate@CDM
WYGIWYS im2latex-100k 37.2 23.9 0.0
I2l-strips im2latex-140k 75.9 65.9 10.3
I2l-nopool im2latex-140k 76.0 66.4 10.4
MathNet im2latexv2 97.2 96.8 83.9
@@
Uni-MuMER Uni-MuMER-Data 99.33 98.71 88.19 93.29
Evaluation on Unnormalized Test Set
Vanilla (baseline) im2latexv2 98.71 97.59 79.85 93.19
Uni-MuMER im2latexv2 98.75 97.85 80.27 93.46
Uni-MuMER Uni-MuMER-Data 98.54 96.72 76.36 93.29